Fermi National Laboratory
 

Accelerator Update for October 12, 2007 - October 15, 2007

 

Friday October 12

The day shift began with the accelerator complex in startup.

Crew Chief Summary: Ground faults from Thursday evening in Tevatron (TeV) and the Main Injector (MI) were found and cleared. The Kautz Road Sub Station (KRSS) power outage was completed on Friday evening. The Collider Detector at Fermilab (CDF) and D-Zero (D0) were searched and secured. The TeV sequencers were tested.

Saturday October 13

The midnight shift began with the accelerator complex in startup.

Crew Chief Summary: Recovered from KRSS outage. Turned on and monitored power supplies in MI. EE support conducted power supply checkouts. Interlock technicians conducted Critical Device Controller (CDC) tests. RF experts worked on the MI and TeV RF. The TeV sector E1 mechanical and cryogenic system work was wrapped up. A MI ramp was established. Operations received permission to begin the TeV cool down and to power the main bus. A sector B0 quadrupole power supply (B0Q5) showed a fault during hipotting. There was TeV separator conditioning. Operations established beam to MI and MiniBooNE. Many people came in to fix problems.

Sunday October 14

The midnight shift began with the accelerator complex in startup.

Crew Chief Summary: TeV experts and EE support investigated the B0Q5 problem – the fault disappeared. MI-8 line studies took place. Pbar studies took place. Operations established beam to the Accumulator. MiniBooNE ran beam to check things out. TeV VCBs were exercised and closed in. A leaky u-tube on the E-1 Dewar was investigated by cryo system techs. They also investigated a TeV sector F4 wet engine problem. MI beam studies took place. Many people came in to fix problems.

Monday October 15

The midnight shift began with the accelerator complex in startup.

Crew Chief Summary: There was a Linac access to repair a vacuum leak on a chopper. TeV CDC testing continued this morning. There will be a TeV access after the CDC testing. Accelerator commissioning/beam startup activities continued.

Machine Summaries for the Weekend

Linac: Delivered beam to Booster. Had a source problem. Vacuum valves closed near the chopper due to a vacuum leak. The system was recovering.

Booster: Circulated beam and sent it to the Main Injector. They had some vacuum problems.

Main Injector: They accelerated beam to 120 and 150 GeV. Sent beam to the Antiproton Source. They conducted aperture scans in MI-40 (it was open), MI-30 (tight vertically), and MI-8, which will need work. They were having problems with bad BPMs and BLMs. They want to smooth the beam in MI-8 and then send beam to the Recycler today. They also want an access later in the week to check out their BPMs and BLMs and to look for possible LCW leaks.

Antiproton Source: They received and circulated beam on the first try. Today they want to conduct aperture scans and to send beam to the Pbar Debuncher. They have problems that will require a tunnel access; they need to fix their momentum system, a bad shunt, a BPM problem, and power supply regulation problem.

Recycler: They are waiting for beam. They reported having had a successful electron cooling study.

SwitchYard 120: F sector separator work was stopped so the technicians could work on another job. They are still in shutdown mode.

NuMI: They are still in shutdown mode.

MiniBooNE: They received beam and have a BPM problem that needs fixing.

Tevatron: They are waiting for the cryo system to get cold, which should be today. They conducted separator conditioning over the weekend. They have permission to power the TeV bus.

D0: Their collision hall was searched and secured on Friday. Their detector was on and they were taking cosmic rays. They reported being ready for beam.

CDF: Their collision hall was searched and secured on Friday. Their detector was on and they were taking cosmic rays. They reported that their Muon Walls on their detector were still open and they would have to have a 1.5-hour access to close them before they could take beam.

FESS: The mussel treatment went well. They reported that the Wilson Hall power outage had been scheduled for Sunday, November 11.

The Plans for Monday

The plans for today are to access the TeV to check the cryo and do a small job in E1. There will be no accesses into F Sector. CDF and D0 will make accesses into their collision halls during the TeV access. Once the TeV to gets cold, experts hope to establish beam into the TeV.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/15/2007    email Fermilab